rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_red
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_red
icon_blue
icon_red
icon_orange
icon_red
icon_red
icon_red
icon_red
icon_red
icon_red
icon_red
icon_green
icon_red
icon_green
icon_green
 

FS#7708 — FS#11594 — Fibre LDN <> AMS

Attached to Project— Network
Incident
the whole network
CLOSED
100%
There is an outage on this fibre. Protection of circuits RBX <> AMS and RBX <> LDN is currently out-of-service.
Date:  Wednesday, 01 October 2014, 17:17PM
Reason for closing:  Done
Comment by OVH - Thursday, 18 September 2014, 10:35AM

We carried out the OTDR measures and found that fault was located on the underwater fibre between London and Amsterdam.

The repair boat is heading towards the cut-off point. It's scheduled to arrive on September 19th.


Comment by OVH - Friday, 19 September 2014, 18:44PM

The boat is on site and has started to remount.


Comment by OVH - Monday, 29 September 2014, 11:58AM

The maintenance team has confirmed that the cable has been fully repaired. Our line is however still down and we don't have control over the Leiston ILA. We will work on the problem with Infinera to remount the line asap and re-establish protections.


Comment by OVH - Wednesday, 01 October 2014, 10:03AM

We're having trouble remounting the link. The Leiston (UK) ILA is offline. A technician is currently on site and we will work on getting service back up with an Infinera engineer.


Comment by OVH - Wednesday, 01 October 2014, 10:05AM

Troubleshoorting highlighted a fault on the OMM card (control card) of the main chassis. A spare is being delivered to replace it. We will need to then reconfigure the new card and then bring the backup on the DB back down to the node.


Comment by OVH - Wednesday, 01 October 2014, 10:06AM

The spare card should arrive around 23:00 CEST.


Comment by OVH - Wednesday, 01 October 2014, 10:09AM

The master chassis is up and we've now got control of the site. Unfortunatley, it seems that the controller card of chassis 2 is also oout-of-service. A spare will therefore be delivered and will arrive on site tomorrow morning. We will start the troubleshooting again tomorrow around 11am with the hope to be able to re-establish the link straightaway.


Comment by OVH - Wednesday, 01 October 2014, 10:12AM

A part of the ramam amps are back in service. The circuits are working one way from Amsterdam > Roubaix. When chassis 2 is back in service, we will be able to remount the other sense.


Comment by OVH - Wednesday, 01 October 2014, 17:17PM

Chassis #2 is back in service. The line has remounted. Protections are active again.